OpenStudy (anonymous):

Give a unique example of each: An absolute value equation with two solutions An absolute value equation with one solution An absolute value equation with no solutions

OpenStudy (anonymous):

|x| = x+5 |x| = 0 |x| = -5

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

yw...:)

jimthompson5910 (jim_thompson5910):

first one |x| = x+5 only has 1 solution

jimthompson5910 (jim_thompson5910):

but something like |x|=3 has two solutions

OpenStudy (anonymous):

oops... sorry... now explanations..

OpenStudy (anonymous):

thank yalll :)

OpenStudy (anonymous):

sorry bout not posting explanations... yw...

OpenStudy (anonymous):

not a prob u help me anyways :)